Snow Valley Ski Resort has been contracting snow removal from its parking lots at a cost of $400/day. A snow-removal machine can be purchased for $27,000. The machine is estimated to have a useful life of 6 years with a zero salvage value at that time. Annual costs for operating and maintaining the equipment are estimated to be $5,250. Determine the break-even value for the number of days per year that snow removal is required in order to justify purchasing the snow-removal machine. MARR is 12%/year.

Break-even value:

days  

Carry all interim calculations to 5 decimal places and then round your final answer up to the nearest day. The tolerance is ±1.
Answer #1

Let the no. of days be n

Annual cost of contracting = 400 * n

Then for breakeven

400 * n = 27000 * (A/P, 12%,6) + 5250

400 * n = 27000 * 0.243226 + 5250 = 11817.10

n = 11817.02 / 400 = 29.54 days ~ 30 days (Nearest day)